Find the absolute maximum and minimum values of the following function over the given interval. If there are multiple points in a single category list the points in increasing order in x value and enter N in any blank that you don't need to use.
1
Expert's answer
2020-04-02T08:39:37-0400

The closed interval method

To find the absolute maximum and minimum values of a continuous function "f" on a closed interval "[a,b]":

1. Find the values of "f" at the critical numbers of "f" in "(a,b)."

2. Find the values of "f" at the endpoints of the interval.

3. The largest of the values from Steps 1 and 2 is the absolute maximum value; the smallest of these values is the absolute minimum value.

"f(x)=6(x^2-1)^3, [-1, 2]"

1. Find the first derivative with respect to "x"


"f'(x)=(6(x^2-1)^3)'=6(3)(x^2-1)^2(2x)=36x(x^2-1)^2"

Find the critical number(s)


"f'(x)=0=>36x(x^2-1)^2=0"

"x_1=-1,x_2=0,x_3=1"

"f(-1)=6((-1)^2-1)^3=0"

"f(0)=6((0)^2-1)^3=-6"

"f(1)=6((1)^2-1)^3=0"


2.

"f(-1)=6((-1)^2-1)^3=0"

"f(2)=6((2)^2-1)^3=162"


3. Absolute maxima


"x=2\ \ \ \ y=162"

Absolute minima


"x=0\ \ \ \ y=-6"